Diagonalise the matrix 𝐴 = [ π‘Ž 𝑏 𝑐 𝑑 ], where π‘Ž + 𝑐 = 𝑏 + 𝑑, by finding a nonsingular matrix 𝑃 and a diagonal matrix 𝐷 such that 𝐴 = 𝑃𝐷𝑃 βˆ’1 .Β 


1
Expert's answer
2021-01-25T14:59:51-0500

"\text{let A =}\begin{bmatrix}
 3 & 4 \\
 2 & 1
\end{bmatrix}\\
\text{to get the eigen values, }\\
t^2-tr(A)t-|A|\\
t^2-4t-5=(t+1)(t-5)\\
\lambda_1=-1, \lambda_2=5\\
\text{to compose the eigen vectors,}\\

\begin{bmatrix}
 3-(-1) & 4 \\
 2 & 1-(-1)
\end{bmatrix}\\
=\begin{bmatrix}
 4 & 4 \\
 2 & 2
\end{bmatrix}\\ 
\text{corresponding to} \\
4x+4y=0\\
\text{the eigen vector is}\\ 
v_1=\begin{vmatrix}
 -1\\
 1
\end{vmatrix}\\
\text{also, }\\

\begin{bmatrix}
 3-(5) & 4 \\
 2 & 1-(5)
\end{bmatrix}\\
=\begin{bmatrix}
 -2 & 4 \\
 2 & -4
\end{bmatrix}\\ 
\text{corresponding to} \\
2x-4y=0, x-2y=0\\
\text{the eigen vector is}\\ 
v_2=\begin{vmatrix}
 2\\
 1
\end{vmatrix}\\
\text{then p}=
\begin{bmatrix}
 -1 & 2 \\
 1 & 1
\end{bmatrix}\\
p^{-1}=\frac{1}{-3}
\begin{bmatrix}
 1 & -2\\
 -1 & -1
\end{bmatrix}\\
\text{setting D to be a diagonal matrix}\\
D=\begin{bmatrix}
 -1 & 0 \\
 0 & 5
\end{bmatrix}\\
A=PDP^{-1}"